# MelCap: A Unified Single-Codebook Neural Codec for High-Fidelity Audio Compression

### v2

@@ Abstract
Neural audio codecs have recently emerged as powerful tools for high-quality and low-bitrate audio compression, leveraging deep generative models to learn latent representations of audio signals. However, existing approaches either rely on a single quantizer that only processes speech domain, or on multiple quantizers that are not well suited for downstream tasks. To address this issue, we propose MelCap, a unified “one-codebook-for-all” neural codec that effectively handles speech, music, and general sound. By decomposing audio reconstruction into two stages, our method preserves more acoustic details than previous single-codebook approaches, while achieving performance comparable to mainstream multi-codebook methods. In the first stage, audio is transformed into mel-spectrograms, which are compressed and quantized into compact single tokens using a 2D tokenizer. A perceptual loss is further applied to mitigate the over-smoothing artifacts observed in spectrogram reconstruction. In the second stage, a Vocoder recovers waveforms from the mel discrete tokens in a single forward pass, enabling real-time decoding. Both objective and subjective evaluations demonstrate that MelCap achieves quality on comparable to state-of-the-art multi-codebook codecs, while retaining the computational simplicity of a single-codebook design, thereby providing an effective representation for downstream tasks. Demos are available at here [^1] https://melcapdemo.github.io/Mel_cap_demo/.
@@ 1 Introduction
Discrete audio tokens generated by neural audio codecs compress continuous audio signals into a compact discrete space while preserving perceptual quality and semantic content Mousavi et al. (2025), enabling reduced storage requirements and faster transmission than continuous embeddings Theis et al. (2017). These tokens serve as an efficient and flexible interface for downstream tasks such as Automatic Speech Recognition (ASR)  Radford et al. (2022)  Hsu et al. (2021), Text-To-Speech generation (TTS)  Peng et al. (2024)  Du et al. (2024b), music generation  Yang et al. (2024), and so on. Audio codecs typically consists of an encoder-quantizer-decoder structure to encode, where the encoder transforms the input waveform into a continuous representation  Langman et al. (2025), The quantizer then maps this continuous representation to a discrete code from a codebook. Finally, the decoder reconstructs the original waveform from the selected code Agustsson et al. (2017). Compression is achieved when the number of bits used to represent the code is smaller than that required for the original audio signal Yang et al. (2020).
@@
Based on the number of quantizers used, quantization method of codecs can be broadly categorized into two types: multiple stage vector quantization  Juang and Gray (1982) and single vector quantization (SVQ). Audio codecs, such as SoundStream  Zeghidour et al. (2021), most commonly use residual vector quantizer (RVQ) for quantization. With iterative residual refinement, the multiple stage vector quantizer can decrease loss of information. Multi-codebook codecs depend on multi-sequence prediction, which reduces efficiency and robustness  Li et al. (2024). Single vector quantization is simpler and particularly useful for downstream generation tasks such as acoustic language models Ye et al. (2025). Recent work such as WavTokenizer  Ye et al. (2024) investigates speech compression using a single codebook. However, existing single quantizer approaches do not take into account more complicated audio signals such as music and environmental sounds.
@@
Neural audio codecs can also be categorized into two types based on the representation they compress: waveform-based and spectral-based approaches. In waveform-based neural codecs, waveforms are directly passed to the encoder. Waveform tokens are typically learned using encoder–decoder architecture trained to reconstruct the waveform Mousavi et al. (2025). EnCodec  Défossez et al. (2022) extends this architecture with a multi-scale STFT discriminator, which help reduce artifacts and produces high-quality samples. DAC  Kumar et al. (2023) improves this framework by introducing multiscale mel reconstruction loss, which better captures details and thus improves audio quality. SNAC  Siuzdak et al. (2024) extends Residual Vector Quantization (RVQ) to multiple temporal resolutions, resulting in more efficient compression.
However, waveform-based approaches still require large model capacity and a greater number of quantizers to capture fine frequency detail accurately seen in Figure  1, which is incompatible with our single-codebook objective. Spectral-based approaches solve this problem by transforming the waveform into the spectral domain, which provides a more effecient representation and allows the model to better capture fine-grained frequency details. Recent works such as APCodec  Ai et al. (2024) jointly models amplitude and phase spectra with residual vector quantization and GAN-based training, enabling high-quality 48 kHz audio reconstruction. However, the instability of Generative Adversarial Network (GAN) training hinder the model’s capacity to faithfully reconstruct the input audio, particularly subtle frequency details and original phase Wu et al. (2024).
To address the aforementioned challenge, this paper proposed a novel audio codec based on mel-spectrogram, which is a compact representation that can compresses complex audio signals into a single codebook. There are three main contributions of our method. First, this codec incorporates perceptual loss into mel-spectrogram reconstruction to alleviate the over-smoothing problem, and further relates it to the feature matching loss used in traditional GAN-based codecs. Second, we use a two-stage training framework to train Vector Quantized-Variational AutoEncoder (VQ-VAE) and GAN-based Vocoder separately, which leads to better GAN training stability and audio quality. Third, this codec aims to encode high-sampling-rate audio (e.g., 44 kHz) using a single quantizer layer, thus meeting the requirements of downstream generation tasks.
@@ 2 Related Work
2.1 2D Transformer Tokenizer
Discrete audio tokenizers often compress audio into a latent one-dimensional representation, which is then quantized into a sequence of discrete tokens. This process is accomplished by compressing the audio along the temporal dimension, so it is defined as a 1D tokenizer. A 2D tokenizer compresses audio in both time and frequency dimensions (typically operating on spectrograms), and then transforms the compressed representation into a sequence of discrete tokens. Many works on 2D tokenizers have been explored in the image domain Yu et al. (2024). VQ-VAE  Yang et al. (2020) first introduced vector quantization in the latent space of VAEs to map images, audio, and video into discrete values. Vector Quantized Generative Adversarial Network (VQGAN)  Esser et al. (2021) extended this by adding perceptual and adversarial losses to better capture detail infermation. ViT-GAN further replaced convolutions with ViT Transformers  Dosovitskiy et al. (2021). In this paper, we assume that the latent space of audio should preserve a 2D structure, maintaining an explicit alignment between time and frequency. Building on the structure of powerful 2D Transformer tokenizers  NVIDIA et al. (2025), we fully explore compact 2D representations for audio.
@@ 2.2 Vocoder
Neural vocoders are neural network models that converts intermediate representations, such as mel-spectrograms, into high-fidelity audio Jiao et al. (2021). Autoregressive models had long been the best-performing vocoders. WaveNet van den Oord et al. (2016), for instance, uses the mel-spectrogram as a local condition. However, its requirement for sequential (sample-by-sample) generation limits streaming efficiency. GAN-based models are capable of generating speech from mel-spectrogram efficiently Kong et al. (2020). Since low latency is a key property for a good codec, we build our model on Vocos Siuzdak (2024). Vocos is a fast neural GAN-based vocoder designed to reconstruct audio from mel-spectrogram through inverse Fourier transform.
@@ 3 Methods
GAN-based end-to-end codecs require dedicated discriminators and multiple codebooks to improve the waveform details, high-frequency components, and phase synchronization, but training these discriminators is time-consuming and convergence is often slow Wu et al. (2023). Consequently, we propose a two-stage codec, where the first stage focuses on mel-spectrogram reconstruction with metric losses, and the second stage incorporates a discriminator to recover high-fidelity waveform from mel discrete tokens. This architecture significantly improves training efficiency, enabling our second-stage model to converge within only 50 epochs.
3.1 Log-Mel Spectrogram
@@
Figure 3: Training paradigm of MelCap.
3.2 First Stage: Mel-Spectralgram Reconstruction
@@
where $\mathrm{VGG}_{l}(\cdot)\in\mathbb{R}^{T\times M\times C}$ denotes the feature maps extracted from the $l$-th layer of a pre-trained VGG-19 network, $L$ is the number of layers considered, and $\alpha_{l}$ is the weight assigned to the $l$-th layer. To further enhance fine details, we fine-tune the tokenizer using a Gram-matrix loss  Gatys et al. (2016), which emphasizes sharper structures.
$\mathcal{L}_{\text{Gram}}=\frac{1}{L}\sum_{l=1}^{L}\sum_{t}\alpha_{l}\left\lVert\mathrm{GM}_{l}(\hat{S})-\mathrm{GM}_{l}(S)\right\rVert_{1},$ (4)
@@ 3.3.3 Training Objectives
Reconstruction Loss. Reconstruction loss refers to the L1 distance between the mel-scaled magnitude spectrograms of the ground-truth waveform and the generated waveform  Kong et al. (2020). Unlike  Yang et al. (2023) that uses 80 mel-spectrogram bins, our setup constrains the number of bins to be no smaller than the mel-spectrogram resolution defined in the first stage, which is 96 mel-spectrogram bins for music and environment sound. This ensures consistency between the first stage and second stage, preventing the loss of high-frequency details when training second stage.
Feature Matching Loss. Feature matching loss measures the learned similarity between a real and generated sample via discriminator features Larsen et al. (2016),  Kumar et al. (2019) Hifi-GAN  Kong et al. (2020) first used it as an additional loss to train the generator of vocoder. In our case, feature matching is used to reduce over-smoothing, serving a similar role as the VGG loss applied in the first stage. Unlike the original setting in Hifi-GAN, where the feature matching loss weight is 2, we increase it to 5.
@@ 4 Experiments
DataSets. The model is trained on the AudioSet dataset, using the entire balanced training subset (bal train), along with the HQ-Conversations dataset Magic Data (2024). The AudioSet covers a wide range of sounds, including human and animal vocalizations, musical instruments and genres, as well as common everyday environmental noises. We evaluate the in-domain reconstruction performance of the codec on the AudioSet test set. We evaluate the out-of-domain reconstruction performance of the codec on a subset of the music genre dataset  Lewtun (2022), which consists of 1,269 randomly selected 10-second clips. All audio files are kept at their original sampling rate of 44 kHz. For each audio sample, mel-scaled spectrograms are computed with the following parameters: FFT size $n_{fft}$=1024, hop size $hop_{n}$ =256, and 96 Mel bins.
@@
Baseline Methods. Our proposed model is compared against DAC  Kumar et al. (2023), SNAC Siuzdak et al. (2024), Spectral Codec  Langman et al. (2025), NVIDIA NeMo Audio Codec and WavTokenizer  Ji et al. (2025). For all baselines, we use the officially released pretrained checkpoints— the 24kHz version for WavTokenizer and 44 kHz versions for other methods, which are publicly available online. For a fair comparison, we additionally trained a DAC model with a smaller number of quantizers (four codebooks) to match the bitrate of our proposed method.
@@ 4.1 Evaluation
We evaluate our models using four primary objective metrics, VISQOL, LSD, Mel Distance, STFT Distance, and two additional reference metrics, UTMOS and V/UV F1. We also performed subjective evaluation with MUSHRA. The primary metrics assess spectral and perceptual fidelity, while the reference metrics are included for informational purposes, as they are designed for speech and may be less reliable in general audio tests.
@@
UTMOS. UTMOS is an automatic mean opinion score (MOS) prediction system that estimates perceptual audio quality and correlates highly with human judgments at sampling rate 16k. However, since we focus on high-frequency details, this 16kHz sampling rate makes UTMOS less suitable for our evaluation.
V/UV F1. F1 measures the classification accuracy of voiced and unvoiced segments. Since AudioSet contains diverse sound categories beyond speech, this metric—originally designed for speech—is only indicative in our setting.
@@ 4.2 Ablation Experiment Result for Mel-spectrogram Reconstruction
To investigate the impact of different loss functions used in first stage on the perceptual quality of the generated audio, we conduct an ablation study using a fixed pretrained vocoder. Specifically, we compare three training settings: (1) using only the reconstruction loss, (2) using reconstruction loss combined with VGG loss, and (3) using reconstruction loss, VGG loss, and an additional Gram matrix (GM) loss. This study allows us to analyze how each component contributes to perceptual fidelity.
Table 1: Comparison of different loss terms used in the first stage. MAE denotes the element-wise L1 difference between the input and output mel-spectrograms after the first stage.
@@
STFT
Dis
$\downarrow$
Mel
Dis
$\downarrow$
@@
Reconstruction Loss 0.26 4.36 0.67 1.68 0.48 0.58 1.31
+ VGG loss 0.31 4.26 0.64 1.68 0.50 0.63 1.31
+ GM loss 0.41 4.24 0.67 1.73 0.58 0.61 1.31
Table 1 compares different loss terms used in the first stage. While adding GM loss reduces the over-smoothing of the reconstructed spectrogram, it introduces artifacts, leading to worse overall metrics. Therefore, we choose not to use GM loss in our final model.

### v3

# MelTok:2D Tokenization for Single-Codebook Audio Compression

@@ Abstract
Large Audio Language Models (LALMs) have emerged with strong performance across diverse audio understanding tasks and can be further enhanced by neural audio codecs. Transitioning from multi-layer residual vector quantizers to a single-layer quantizer has been shown to facilitate more efficient downstream language models decoding. However, the ability of a single codebook to capture fine-grained acoustic details remains limited, as the frequency-variant nature of 1D tokenizers leads to redundancy. To address this issue, we propose MelTok, a two-dimensional (2D) tokenizer that effectively compresses acoustic details of 44.1 KHz audio into a single codebook. The tokenizer encodes audio into a more compact representation than one-dimensional tokenizers. Furthermore, to recover audio from mel-spectrogram tokens, we propose a token-based vocoder. Both objective and subjective evaluations demonstrate that MelTok achieves quality comparable to multi-codebook codecs and outperforms existing state-of-the-art neural codecs with a single codebook on high-fidelity audio reconstruction. By preserving acoustic details, MelTok offers a strong representation for downstream understanding tasks.
Figure 1: Mel-spectrogram comparison of original and reconstructed waveforms produced by different codecs. MelTok accurately reconstructs the high-frequency details.
@@ 1 Introduction
Discrete audio tokens generated by neural audio codecs compress continuous audio signals into a compact discrete space while preserving perceptual quality and semantic content Mousavi et al. (2025). Discrete audio tokens require less storage and achieve faster transmission than continuous embeddings Theis et al. (2017). These tokens serve as an efficient and flexible interface for downstream audio understanding tasks such as Automatic Speech Recognition (ASR)  Radford et al. (2022a)  Hsu et al. (2021), sound event detection Cakir et al. (2015), music genre classification  Bahuleyan (2018), and so on. Audio codecs typically consists of an encoder-quantizer-decoder structure to encode. The encoder transforms the input waveform into a continuous representation  Langman et al. (2025), The quantizer then maps this continuous representation to a discrete code from a codebook. Finally, the decoder reconstructs the original waveform from the selected code Agustsson et al. (2017). Compression is achieved when the number of bits used to represent the code is smaller than that required for the original audio signal Yang et al. (2020).
@@
Based on the number of quantizers used, quantization method of codecs can be broadly categorized into two types: multiple stage vector quantization  Juang and Gray (1982) and single vector quantization (SVQ). Full-Bandwidth Audio Codecs, such as DAC  Kumar et al. (2023), most commonly use residual vector quantizer (RVQ) for quantization. With iterative residual refinement, the multiple stage vector quantizer can compress 44.1 KHz audio with minimal loss in quality Kumar et al. (2023). However, multi-codebook codecs depend on multi-sequence prediction, which reduces efficiency and robustness  Li et al. (2024). Single vector quantizers has emerged as simpler and particularly useful tools for downstream generation tasks such as acoustic language models Ye et al. (2025). Recent work such as WavTokenizer  Ye et al. (2024) and UniCodec  Jiang et al. (2025) investigates compression in different domains using a single codebook. However, existing single quantizer approaches do not take “full bandwidth” rates of 44.1 or 48 kHz into account. This limitation is critical, as high-resolution acoustic detail is often required for good performance in general downstream tasks. Valin et al. (2016).
@@
Neural audio codecs can also be categorized into two types based on the representation they compress: waveform-based and spectral-based approaches. In these waveform-based codecs, 1d tokenizers are naturally used to compress one-dimensional time-domain data Mousavi et al. (2025). DAC  Kumar et al. (2023) introduces multiscale mel reconstruction loss into this framework, which better captures details and thus improves audio quality. SNAC  Siuzdak et al. (2024) extends Residual Vector Quantization (RVQ) to multiple temporal resolutions, resulting in more efficient compression. In these waveform-based codecs, 1d tokenizers are naturally used to compress one-dimensional time-domain data.
However, high-frequency modeling is still challenging for these waveform-based models especially at high sampling rates with fewer quantizers seen in Figure  2, which is incompatible with our single-codebook objective. Spectral-based approaches solve this problem by transforming the waveform into the two-dimensional frequency domain features, which provides a more effecient representation and allows the model to better capture fine-grained frequency details Guo et al. (2025). Recent works such as Spectral CodecLangman et al. (2025) compress the mel-spectrogram and reconstruct the time-domain audio signal, which typically rely on 1D convolution kernels operating along the temporal axis. 2D convolution kernels are first used in FunCodec Du et al. (2023) after transforming waveform into spectrum, followed by a residual vector quantization (RVQ) module. While most recent codecs adopt a 1D tokenizer, they often suffer from inherent limitations: the resulting representations are not sufficiently compact. And they imperfectly model high-frequency content through one codebook, leading to audio that are clearly distinguishable from originals  Défossez et al. (2022). Our contributions are as follows:
Compression of 44.1 KHz audio in one single codebook: This codec aims to encode high-resolution audio using a single quantizer layer with minimal loss of high-frequency content, which is crucial for downstream audio understanding tasks.
Perceptual loss in mel-spectrogram reconstruction: This codec incorporates perceptual loss into mel-spectrogram reconstruction to alleviate the over-smoothing problem, and further relates it to the feature matching loss used in traditional GAN-based codecs.
A ”frequency-invariance” representation for audio: We identify a critical issue in existing single quantizer codecs which can not reconstruct full bandwidth audio due to frequency-variance redundancy. Unlike classical 1D tokenizers, which treat frequency content sequentially and are thus constrained by the assumption of frequency variance, MelTok utilizes a 2D tokenizer.
Token-based vocoder:We use a two-stage training framework to train 2D tokenizer and token-based Vocoder separately, which leads to better GAN training stability and audio quality.
@@ 2 Related Work
2.1 2D Tokenizer
Existing discrete audio tokenizers compress the waveform or spetrogram along the temporal dimension through Conv1D, so it is defined as a 1D tokenizer. A 2D tokenizer compresses 2 dimentional data in both time and frequency dimensions, and then transforms the compressed representation into a sequence of discrete tokens. Many works on 2D tokenizers have been explored in the image domain Yu et al. (2024). VQ-VAE  Yang et al. (2020) first introduced vector quantization in the latent space of VAEs to map images, audio, and video into discrete values. Vector Quantized Generative Adversarial Network (VQGAN)  Esser et al. (2021) extended this by adding perceptual and adversarial losses to better capture detail infermation. ViT-GAN further replaced convolutions with ViT Transformers  Dosovitskiy et al. (2021).
Some work in audio has also explored the usage of 2D convolutional layers. For example, PANNs  Kong et al. (2020b) employ 2D CNNs such as CNN14 to capture time–frequency invariant patterns on Wavegram features, which helps the model improves Audio Pattern Recognition performance. To explicitly capture harmonic structure, a cross-frequency structure, Ulicny et al. (2022) proposed a 2D convolution Harmonic Convolution. Inspired by the success of 2D kernels used in capturing audio structures, we aim to leverage this invariance to derive a more compact discrete representation for high-fidelity neural audio codec. Building on the structure of powerful 2D Transformer tokenizers  NVIDIA et al. (2025), we fully explore compact 2D representations for audio.
@@ 2.2 Vocoder
Neural vocoders are neural network models that converts intermediate representations, such as mel-spectrograms, into high-fidelity audio Jiao et al. (2021). Autoregressive models had long been the best-performing vocoders. WaveNet van den Oord et al. (2016), for instance, uses the mel-spectrogram as a local condition. However, its requirement for sequential (sample-by-sample) generation limits streaming efficiency. GAN-based models are capable of generating speech from mel-spectrogram efficiently Kong et al. (2020a). Since low latency is a key property for a good codec, we build our model on Vocos Siuzdak (2024). Vocos is a fast neural GAN-based vocoder designed to reconstruct audio from mel-spectrogram through inverse Fourier transform.
@@ 3 Methods
To achieve faster convergence, we propose a two-stage codec, where the first stage focuses on mel-spectrogram reconstruction with metric losses, and the second stage incorporates a discriminator to recover high-fidelity waveform from mel discrete tokens. This architecture significantly improves training efficiency, enabling our second-stage model to converge within only 50 epochs.
@@
3.1 2D Tokenizer VS 1D Tokenizer
Figure 3: t-SNE visualization of latent embeddings produced by the 1D and 2D tokenizers. Each point represents the latent embedding of a single time frame, and each color denotes a different frequency band.
Compared with conventional 1D tokenizers, our proposed 2D tokenizer provides a more compact representation for single-codebook audio compression. A 1D tokenizer processes only the temporal dimension, forcing the model to encode all frequency variation into a single token sequence. In contrast, our 2D tokenizer maintains the inherent time–frequency structure, decreasing redundancy caused by frequency variation. Furthermore, patchifying the mel-spectrogram enables the two-dimensional tokenizer to generate frequency-specific embeddings. The detail of patchify can be seen in  A.7. This leads to disentangled representations in which high-, mid-, and low-frequency structures are clearly separated in the latent space, seen in  3. These findings suggest that 2D tokenization provides a more suitable representation than 1D tokenization for single-codebook audio codecs. The encoder details are given in  7.
Figure 5: Training paradigm of MelTok.
3.2 First Stage: Mel-Spectralgram Reconstruction via 2D Tokenizer
@@
where $\mathrm{VGG}_{l}(\cdot)\in\mathbb{R}^{T\times M\times C}$ denotes the feature maps extracted from the $l$-th layer of a pre-trained VGG-19 network, $L$ is the number of layers considered, and $\alpha_{l}$ is the weight assigned to the $l$-th layer.
@@ 3.3.3 Training Objectives
Reconstruction Loss. Reconstruction loss refers to the L1 distance between the mel-scaled magnitude spectrograms of the ground-truth waveform and the generated waveform  Kong et al. (2020a). Unlike  Yang et al. (2023) that uses 80 mel-spectrogram bins, our setup constrains the number of bins to be no smaller than the mel-spectrogram resolution defined in the first stage, which is 96 mel-spectrogram bins for music and environment sound. This ensures consistency between the first stage and second stage, preventing the loss of high-frequency details when training second stage.
Feature Matching Loss. Feature matching loss measures the learned similarity between a real and generated sample via discriminator features Larsen et al. (2016),  Kumar et al. (2019) Hifi-GAN  Kong et al. (2020a) first used it as an additional loss to train the generator of vocoder. In our case, feature matching is used to reduce over-smoothing, serving a similar role as the VGG loss applied in the first stage. Unlike the original setting in Hifi-GAN, where the feature matching loss weight is 2, we increase it to 5.
@@ 4 Experiments
DataSets. The model is trained on the AudioSet dataset, using the entire balanced training subset (bal train), musdb18-hq and the HQ-Conversations dataset Magic Data (2024). The AudioSet covers a wide range of sounds, including human and animal vocalizations, musical instruments and genres, as well as common everyday environmental noises. We evaluate the reconstruction performance of the audio domain on the AudioSet evaluation set. We evaluate the reconstruction performance of the music domain on the mixture sound in musdb18hq test set. All audio files are kept at their original sampling rate of 44 kHz. For each audio sample, mel-scaled spectrograms are computed with the following parameters: FFT size $n_{fft}$=1024, hop size $hop_{n}$ =256, and 96 Mel bins.
@@
Baseline Methods. Our proposed model is compared against DAC  Kumar et al. (2023), SNAC Siuzdak et al. (2024), Spectral Codec  Langman et al. (2025), NVIDIA NeMo Audio Codec, UniCodec  Jiang et al. (2025) and WavTokenizer  Ji et al. (2025). For all baselines, we use the officially released pretrained checkpoints— the 24kHz version for WavTokenizer and 44 kHz versions for other methods, which are publicly available online.
@@ 4.1 Evaluation
We evaluate our models using five primary objective metrics, VISQOL, LSD, Mel Distance, STFT Distance, and Mel Cepstral Distortion. We also performed subjective evaluation with MUSHRA. The primary metrics assess spectral and perceptual fidelity.
@@
MCD. Mel Cepstral Distortion (MCD) is a measure of how different two sequences of mel cepstra are. The core idea is that the smaller the MCD between the original and reconstructed mel cepstral sequences, the more accurately the codec preserves the spectral characteristics and perceptual quality of the original audio signal.
For audio dataset, we use Frechet Audio Distance (FAD), which is a distribution-level metric that measures how similar reconstruted audio is to a input corpus by comparing their distributions in a deep embedding space. In our setup, we extract embeddings using a pretrained CLAP model Wu et al. (2024b) with a sample rate of 48 kHz.
For music dataset, we also use another two metrics: We calculate the similarity between the reconstruted vocal part and the original vocal part with a pre-trained wavlm-base-sv  Chen et al. (2022) and use Whisper-tiny  Radford et al. (2022a) to compute the Word Error Rate (WER) of the generated vocal part as the vocal clarity evaluation metric.
@@ 4.2 Ablation Experiment Result for Mel-spectrogram Reconstruction
To investigate the impact of 2D tokenizer used in first stage on the perceptual quality of the generated audio, we conduct an ablation study using a fixed pretrained vocoder. Specifically, we input the reconstructed mel-spectrograms obtained from 1D and 2D tokenizer into a vocoder trained with ground truth mel-spectrograms.
Table 1: Ablation Experiment One: Comparison of 1d tokenizer and 2d tokenizer used in the first stage.
@@
1D Tokenizer 3.56 1.09 11.47 3.97 1.28
2D Tokenizer 4.36 0.67 2.64 1.68 0.48
Table 1 compares different tokenizers used in the first stage. This demonstrates that incorporating using 2d tokenizer can improves the mel-spctrogram reconstruction, which is beneficial for audio reconstruction in the second stage. We provide the encoder architecture details of 1D tokenizer and 2D tokenizer in appendix  8.
Table 2: Ablation Experiment Two: Comparison of different loss terms used in the first stage.
@@
w/o vgg 4.18 0.76 3.82 1.87 0.57
w/ vgg 4.29 0.66 3.74 1.90 0.56
Table 2 compares different loss terms used in the first stage. This demonstrates that incorporating VGG loss effectively mitigates over-smoothing and enhances spectral reconstruction , which is beneficial for training in second stage.